4. Which points on the graph of y = 5 – x2 are closest to the point (0,4)?write clearly please
Answer #1

u Giren y= - 5-x² |(9,5) ز 1 (14) = 5-u 2 (1479 > y = Gorggh 0 3 Johat- 2 (21) 2 17 4 + 3 -2 + 3 4 s 2 -4 -31-2 -1 4 i Å ♡ -4